OpenStudy (anonymous):

solve: 2^(2x+1) - 33(2^x) +16 = 0

OpenStudy (asnaseer):

let \(y=2^x\) and substitute into your equation. you should end up with a quadratic in y. solve that, then find x.

OpenStudy (anonymous):

0_o

OpenStudy (asnaseer):

remember that \(2^{2x+1}=2\times2^{2x}=2\times(2^x)^2\)

OpenStudy (anonymous):

ah!

OpenStudy (asnaseer):

:)

OpenStudy (anonymous):

thank you so much

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

maths is amazing 0_o

OpenStudy (asnaseer):

it's good when you get those "aha!" moments! :D

OpenStudy (anonymous):

:)